OpenStudy (anonymous):

PLEASE HELP!!!!!! (medal will be given to best answer) Each letter of the alphabet is written on a card. The cards are placed into a bag. Find the indicated probability: Three vowels are drawn without replacement. Please explain how you'd solve this problem. Thanks! :)

OpenStudy (anonymous):

\[\frac{5}{26}\times \frac{4}{25}\times \frac{3}{24}\]

OpenStudy (anonymous):

Yay! I got it right! Thank you very much!

OpenStudy (anonymous):

yw